# Mapping the Podcast Ecosystem with the Structured Podcast Research Corpus

### v1

@@ Abstract
Podcasts provide highly diverse content to a massive listener base through a unique on-demand modality. However, limited data has prevented large-scale computational analysis of the podcast ecosystem. To fill this gap, we introduce a massive dataset of over 1.1M podcast transcripts that is largely comprehensive of all English language podcasts available through public RSS feeds from May and June of 2020. This data is not limited to text, but rather includes audio features and speaker turns for a subset of 370K episodes, and speaker role inferences and other metadata for all 1.1M episodes. Using this data, we also conduct a foundational investigation into the content, structure, and responsiveness of this ecosystem. Together, our data and analyses open the door to continued computational research of this popular and impactful medium.
@@ 1 Introduction
Over the past two decades, podcasts have emerged as an important part of the modern media landscape (Aufderheide et al., 2020). With its low barrier to entry, this medium attracts many independent voices, but has also been embraced by professional media and celebrities, collectively supported by billions of dollars in advertising Wirtschafter (2023). However, despite significant evidence highlighting the wide-scale distribution and real-world impact of podcasts (Bottomley, 2015), there is only limited understanding of this medium and even fewer methods for analyzing podcasts, in part due to technical hurdles and limited data availability. In this work, we present the Structured Podcast Research Corpus (SPoRC), the first large-scale dataset of podcasts for NLP and computational social science research, and use it to analyze and document the podcast ecosystem at scale.
Access to large-scale text data has been fundamental to enabling new research directions. Past resources for obtaining data from Reddit (Baumgartner et al., 2020), Twitter (Pfeffer et al., 2023), and Amazon (Ni et al., 2019), have all led to multiple concomitant advances in modeling, applications, and understanding of human behavior. Podcasts, however, introduce unique challenges as a source of data. As an audio medium, the text for podcast episodes is inaccessible in its original form; further, unlike structured media from single platforms, podcasts are distributed by multiple sources and varied metadata exist for each episode. Here, we introduce SPoRC to overcome these challenges.
SPoRC is a thick sample of 1.1M podcast episodes from May–June 2020, each associated with transcripts, hosting platform, inferred host and guest names, and other metadata. Using a highly-parallelized pipeline for the collection and processing of podcasts, we introduce new data and models for extracting hosts and guests from podcast transcripts, enabling studies of the social network. We further use these identities to link specific turns of dialog to their speakers, where possible.
@@
In sum, our paper makes the following three contributions. First, we introduce SPoRC and the computational infrastructure and models needed to create it (§3). Second, we present the first large-scale characterization of the content and structure of the podcast ecosystem; these analyses reveal the topical distribution discussed on podcasts, and the community network structure created by guest co-appearances (§4). Third, we analyze the responsiveness of podcasts, examining the timing and extent of the impact of a major media event (§5), finding varied responses across different segments of the medium. Overall, our results point to new fundamental questions for further research into phenomena such as community identity, information diffusion, and incidental news exposure in podcasts. With the release of SPoRC, these and other lines of inquiry can be quantitatively studied at scale. Our code and data are made available for non-commercial use and can be found at https://github.com/blitt2018/SPoRC_data.
@@ 2 Background
Podcasts are a massively popular, diverse, and impactful medium. In 2023, 42% of Americans aged 12 and older reported having listened to a podcast in the past month, with many listening multiple times per week Pew Research Center (2023b). Podcasts feature long form educational content from leading experts alongside short morning news shows, true crime dramas, life advice, and comedy. These shows also have real-world impact; listeners report shifting their media diet (60%), changing their lifestyle (36%), buying a product (28%), or contributing to a political cause (13%), as a result of their consumption Pew Research Center (2023d). Importantly, the majority of podcast listeners trust information they hear on these shows Pew Research Center (2023d), even as evidence accumulates about podcasts spreading misinformation (Brandt et al., 2023; Wirtschafter, 2021, 2023).
@@
Despite their popularity and reach, there has been surprisingly little large scale research into podcasts, in part due to a lack of data. The largest existing research corpus is a dataset of 200K podcasts (half English; half Portuguese) created by Spotify Clifton et al. (2020). But that dataset is no longer being maintained, and Spotify is no longer granting access to new users. A smaller corpus, with 240 hours of podcasts, annotated with emotion labels, was released in 2019 (Lotfian and Busso, 2019). Another project, PodcastRE, is actively preserving podcast audio, but the focus of that project is primarily archival, and it does not widely distribute data for research Morris et al. (2019).
As with comparable resources for other types of data source, such as Congressional Speeches (Gentzkow et al., 2019) or Reddit (Baumgartner et al., 2020), our dataset enables new research into podcasts, with speaker-labeled content over a subset that is dense enough for the study of communities, yet long enough to study temporal trends. Compared to other media, our dataset capture long form conversations, augmented with audio features that give insight into how each speaker is communicating, with speakers appearing in many venues, connected via content, guests, and other mechanisms. Towards this end, we carry out two initial high-level analyses (§4 and §5) that make use of multiple dimensions of this data to study the structure and responsiveness of the podcast ecosystem.
3 Building our podcast dataset
To create a comprehensive and open dataset for research into the podcast ecosystem, or for use in studying related social scientific and linguistic questions, we set out to collect and process all extant podcast content from a fixed slice of time. Because our goal is not to redistribute audio files, we only preserve and share transformed versions of original podcast audio, as described below.111For the purposes of this work, we consider podcasts to be any serially-released audio recordings distributed as mp3 files via RSS feeds; note that this excludes video podcasts that are only distributed on platforms such as YouTube. Here, we use the term “podcast” and “show” synonymously, where each podcast consists of or more episodes.
@@ 3.1 Initial data collection
To bound the scope of the initial dataset, we chose the months of May and June, 2020, and attempt to collect all podcast audio files from that time that are still available from RSS feeds.222Choosing a time period from the recent past allows us to understand the data with some context; notable events during this time period include the murder of George Floyd and the ensuing protests, the start of Operation Warp Speed, and US deaths from COVID-19 passing 100K people. To identify RSS feeds, we start with data from Podcast Index, which includes information on over 4 million different shows, complete with podcast-level metadata.333https://podcastindex.org/ Using the RSS feeds provided by Podcast Index, we identify 273K English-language shows that released episodes during May or June, 2020. From these feeds, we are able to download both audio files and episode-level metadata for a total of 1.3M episodes from 247K different shows.
@@ 3.2 Transcription
To extract the words that are spoken, we begin by transcribing each audio file to a plain text format. For this, we use Whisper, a publicly available automated speech recognition (ASR) system (Radford et al., 2022). Specifically, we use the “whisper-base” model, which represents a compromise between quality and speed.444https://huggingface.co/openai/whisper-base Whisper also provides approximate timestamps at the word-level, as well as tags such as “(laughing)” and “[MUSIC]” to mark non-speech audio segments. Although prior work has identified some important issues with Whisper, including a propensity to hallucinate (Koenecke et al., 2024), we find that this is largely not an issue for our data, and take additional steps to mitigate this potential problem (see Appendix A for details). After filtering out failed transcriptions, we are left with 1.1 million episode transcripts, comprising 6.6 billion words.
@@ 3.3 Prosodic Feature Extraction
In addition to textual content, podcast audio contains rich information about how a speaker is communicating. To obtain this additional information, we use the openSMILE toolkit (Eyben et al., 2010) focusing on a subset of the eGEMAPS feature set (Eyben et al., 2015). In particular, we extract the fundamental frequency (F0), the first formant (F1), the first four Mel Frequency Cepstral Coefficients (MFCCs 1-4). F0 is related to vocal pitch, and captures the lowest frequency of oscillations of a voiced sound. In addition to varying across speakers, speakers manipulate this frequency in speaking, such as to mark emphasis or questions. Formants capture additional resonances or harmonics, with F1 being commonly associated with vowel pronunciation. It has been further employed to describe variation in speakers’ dialect, sex, and age Hagiwara (1997); Kent and Vorperian (2018). Finally, MFCCs are features designed to capture the short-term power spectrum of sounds, and can be used in concert with other features to measure general vocal characteristics for a number of downstream tasks such as emotion recognition Luengo et al. (2005). Because openSMILE measures these features at a high frequency, with correspondingly large storage requirements, we collapse these to token-level metrics, keeping the mean of each feature over the course of each word.
@@ 3.4 Identifying Speakers
A key limitation of Whisper is that it does not distinguish between speakers. In order to measure turn-level information from podcast conversations, we attempt to match speaking turns to speakers.
Speaker Diarization
@@
As a first step, we use pyannote (Bredin, 2023) to split audio files into individual speaker turns, using a process known as diarization.555Due to computational limitations and the speed of diarization, we only diarize 370K episodes in our initial dataset release, sampled randomly from the full set of transcribed episodes, but will update this in subsequent data releases. pyannote determines the number of distinct speakers in a conversation, matching these speakers to generic labels (e.g., speaker1, speaker2, etc.). Each audio segment is then assigned to one of these speakers, which allows us to map each token from transcripts to a corresponding speaker turn, using audio timing information. After excluding speakers that make up less than 5% of the total speaking time in an episode, we find that 37% of episodes in our diarized data have only a single speaker, with 39% having two speakers, and the rest having three or more. For additional details on diarization, please refer to Appendix B.
Speaker Labeling
@@
Unfortunately, mapping generic speaker IDs to named speakers or speaker roles is difficult, as this information is not typically included in episode metadata in a structured format. As a step towards this, we begin by attempting to determine the names and voices of the host(s) and guest(s) for each episode.
@@
Conveniently, hosts and guests are often mentioned by name in episode descriptions and transcripts (e.g., “I am your host …”), but not necessarily in a systematic way.666In addition, some hosts never introduce themselves within a given episode, or provide only their first name, and thus remain effectively anonymous in our data. We develop a pipeline to first identify candidate names in text, and then classify each of these names as Host, Guest, or Neither, with the third category representing people mentioned in but not appearing in the episode. To make these assignments, we rely on the direct and indirect linguistic cues surrounding mentions of a speaker’s name that indicate their role. We first use spaCy Honnibal et al. (2020) to identify all named people that occur in an episode description, or in the first 350 words of the transcript, filtering this list to keep only mentions that include a first and last name.
@@
To generate training data for this task, we collect three human judgements from Profilic annotators on a subset of 2,000 candidate entities in context. Annotators label one target entity at a time, and see the podcast description, episode description, and 300 words of the transcript, and provide a judgement of Host, Guest, or Neither. Despite some ambiguous cases, annotators had relatively high chance-corrected agreement on this task (Krippendorff’s =0.77). Final labels were obtained by aggregating annotations using MACE Hovy et al. (2013), resulting in labels for 858 Hosts, 639 Guests, and 503 Neither. Using these annotations, we fine tune a RoBERTa model to classify a named entity based on its mean-pooled embeddings, which achieves 0.87 cross validation accuracy and 0.88 accuracy on a held out test set. Finally, for diarized podcasts with an identified Host, we heuristically map that person to the first voice which says the host’s name. For additional details on speaker labeling, see Appendix C.
@@ Results
Using our role annotation model, we identify participants in all episodes; over 550K episodes (49%) of the dataset have at least one identified Host or Guest, including 386K Host names and 535K Guest names. Based on these inferred participants, we find that the plurality of podcasts have a single Host, followed by a single Host and single Guest (see Figure 5 in the Appendix). During manual review, we found a few notable false positives, such as Ryan Reynolds, who appears in a common podcast ad in which he is labeled as a Host. Nevertheless, we verify that some guests truly are prolific, such as Matt Ridley, who appears as a guest on 21 episodes in our data. Further refinements could improve on our host and guest identification by combining information across episodes and using external resources, but these labels nevertheless permit further investigation into how guests connect shows across categories (§4).
@@ 3.5 Summary of Our Dataset
Altogether, our data includes rich multimodal information spanning the episode and speaker turn levels. At the episode level, we provide transcript text alongside category, duration, publication date, and inferred host and guest information. At the turn level, we provide transcripts split into segments of speech corresponding to different speakers, start and end timestamps, averaged audio features for each segment, and their inferred host and guest roles where possible. We release episode-level data for the full set of 1.1 million podcast episodes and speaker turn data for the subset of 370K diarized episodes (see Appendix E for details).

### v2

@@ Abstract
Podcasts provide highly diverse content to a massive listener base through a unique on-demand modality. However, limited data has prevented large-scale computational analysis of the podcast ecosystem. To fill this gap, we introduce a massive dataset of over 1.1M podcast transcripts that is largely comprehensive of all English language podcasts available through public RSS feeds from May and June of 2020. This data is not limited to text, but includes metadata, inferred speaker roles, and audio features and speaker turns for a subset of 370K episodes. Using this data, we conduct a foundational investigation into the content, structure, and responsiveness of this ecosystem. Together, our data and analyses open the door to continued computational research of this popular and impactful medium.
@@ 1 Introduction
Over the past few decades, podcasts have emerged as an important part of the modern media landscape (Pew Research Center, 2023b). Despite their growing importance, most published research on podcasts has used small samples and/or non-computational methods to study topics like educational content (Quintana and Heathers, 2021), misinformation (Wirtschafter, 2023), and listener preferences (Moe, 2022), while largely neglecting distinguishing aspects of this medium, such as its audio characteristics and social network.
To enable research into media, communication, and the podcast ecosystem, we introduce the Structured Podcast Research Corpus (SPoRC), the first large-scale open dataset of podcasts providing a nearly comprehensive slice of publicly available data to support computational social science research. In this paper, we present SPoRC and our process for building it, along with preliminary analyses of the dataset, exploring the breadth of topical content, network linkage across shows, and responsiveness to real-world news events.
Given that podcasts are already a key channel for people to consume entertainment and information (Tobin and Guadagno, 2022; Pew Research Center, 2023a), one cannot fully understand the modern media landscape without considering long-form audio content. However, in comparison to domains like news or social media, there is currently a dearth of available datasets for studying this medium at scale. By creating a large-scale corpus that surfaces both text and audio aspects of podcasts, enhanced with additional metadata, we aim to facilitate the kinds of insights that have come from similarly scaled public research datasets for Twitter (Pfeffer et al., 2023), Reddit (Baumgartner et al., 2020), Amazon (Ni et al., 2019), and other domains.
While the long term evolution of podcasts is also of great interest, the initial SPoRC dataset is a thick slice of shows, representing every English language episode we were able to retrieve from the months of May and June 2020—a total of 1.1M episodes with transcripts and basic metadata. Diarization and speaker identification have also been included for a subset of episodes. This sample thus provides a foundation for future understanding, including the long tail of podcast content.
@@
In the rest of the paper, we first introduce SPoRC and the methods used to create it (§3). We then characterize this snapshot of the podcast ecosystem in terms of content and structure, revealing the distribution of topics discussed on podcasts, and the community network structure created by guest co-appearances (§4). Finally, we analyze the responsiveness of podcasts, examining the timing and extent of the impact of a major media event (§5), finding varied responses across different segments of the medium. Overall, our results point to new questions for further research into community identity, information diffusion, and incidental news exposure in podcasts, which our release of SPoRC will enable. Our data and code are made available for non-commercial use and can be found at https://github.com/blitt2018/SPoRC_data and https://huggingface.co/datasets/blitt/SPoRC.
@@ 2 Background
Podcasts are a massively popular, diverse, and impactful medium. In 2023, 42% of Americans aged 12 and older reported having listened to a podcast in the past month, with many listening multiple times per week Pew Research Center (2023a). Podcasts feature long-form educational content alongside short morning news shows, true crime dramas, life advice, and comedy. These shows also have real-world impact; listeners report shifting their media diet (60%), changing their lifestyle (36%), buying a product (28%), or contributing to a political cause (13%), as a result of their consumption Pew Research Center (2023b). Importantly, the majority of podcast listeners trust information they hear on these shows Pew Research Center (2023b), even as evidence accumulates about podcasts spreading misinformation (Brandt et al., 2023; Wirtschafter, 2021, 2023).
@@
Despite their popularity and reach, there has been surprisingly little large-scale academic research into podcasts, in part due to a lack of data. The largest existing research corpus is a dataset of 200K episodes (half in English; half in Portuguese) released by Spotify Clifton et al. (2020). But that dataset is no longer being maintained, and Spotify is no longer granting access to new users. A smaller corpus, with 240 hours of podcasts, annotated with emotion labels, was released in 2019 (Lotfian and Busso, 2019). Another project, PodcastRE, is actively preserving podcast audio, but the focus of that project is primarily archival, and it does not widely distribute data for research Morris et al. (2019). Finally, the Brookings Institute has assembled and made available the Popular Political Podcasts Dataset (PPPD), a corpus of around 120 prominent political podcast, which is actively being maintained (Brandt et al., 2023).
As with comparable resources for other types of data sources, such as Congressional Speeches (Gentzkow et al., 2019) or Reddit (Baumgartner et al., 2020), our dataset enables new research into podcasts, with speaker-labeled content over a subset that is dense enough for the study of communities, yet long enough to study temporal trends. Compared to other media, our dataset captures long-form conversations, augmented with audio features that give insight into how each speaker is communicating, with speakers appearing in many venues, connected via content, guests, and other mechanisms. Towards this end, we carry out two initial high-level analyses (§4 and §5) that make use of multiple dimensions of this data to study the structure and responsiveness of the podcast ecosystem. We also intend that the pipeline we have created (§3) will be useful to others in curating additional podcast datasets.
3 Building a Massive Podcast Dataset
To create a comprehensive and open dataset for research into the podcast ecosystem, or for use in studying related social scientific and linguistic questions, we set out to collect and process all extant podcast content available from RSS feeds, covering a fixed slice of time.111For the purposes of this work, we consider podcasts to be any serially-released audio recordings distributed as mp3 files via RSS feeds; note that this excludes video podcasts that are only distributed on platforms such as YouTube. Here, we use the term “podcast” and “show” synonymously, where each podcast consists of one or more episodes. For additional discussion of how to define the medium, see McGregor (2022).,222RSS feeds are the standardized format used to structure podcast metadata, allowing them to be streamed across a variety of platforms. Although the vast majority of podcasts are freely available and portable in this way, a subset have exclusive deals with specific hosting platforms preventing public access to their feeds. Although long term dynamics of podcasts are also of great interest, SPoRC provides a foundation for future exploration, and can easily be extended by others using code that we will release. Because our goal is not to redistribute audio files, we only preserve and share transformed versions of original podcast audio, as described below.
@@ 3.1 Initial Data Collection
To bound the scope of the initial dataset, we chose the months of May and June, 2020, and attempt to collect all podcast audio files from that time that are still available from RSS feeds.333Choosing a time period from the recent past allows us to understand the data with some context; notable events during this time period include the murder of George Floyd and the ensuing protests, the start of Operation Warp Speed, and US deaths from COVID-19 passing 100K people. To identify RSS feeds, we start with data from Podcast Index, a public database which includes information on over 4 million different shows, complete with podcast-level metadata.444https://podcastindex.org/ Using the RSS feeds provided by Podcast Index, we identify 273K English-language shows that released episodes during May or June, 2020. From these feeds, we are able to download both audio files and episode-level metadata for a total of 1.3M episodes from 247K different shows.
@@ 3.2 Transcription
To extract the words that are spoken, we begin by transcribing each audio file to a plain text format. For this, we use Whisper, a publicly available automated speech recognition (ASR) system (Radford et al., 2022). Specifically, we use the whisper-base.en model, which represents a compromise between quality and speed.555https://huggingface.co/openai/whisper-base.en Whisper also provides approximate timestamps at the word-level, as well as tags such as “(laughing)” and “[MUSIC]” to mark non-speech audio segments.
Although prior work has identified some important issues with Whisper, including a propensity to hallucinate continuations (Koenecke et al., 2024), we find that this is not a major issue for our data. Rather, the biggest source of errors is that the model will sometimes repeat short phrases many times in a row during segments featuring music, silence, or non-English speech. To maintain high quality transcripts, we use a simple n-gram filter to identify and remove episodes that have a high proportion of this sort of repetition. Regardless, overall transcription quality remains high: validation against a sample of professional podcast transcripts reveals word error rates of less than 10% on average. Moreover, closer inspection reveals that more than half of these errors are due to the way the professional transcripts have been edited for fluency. (See Appendix A for details). After filter, we are left with 1.1 million episode transcripts, comprising 6.6 billion words.
@@ 3.3 Prosodic Feature Extraction
In addition to textual content, podcast audio contains rich information about how a speaker is communicating. To obtain this additional information, we use the openSMILE toolkit (Eyben et al., 2010), which is widely used for audio feature extraction in scientific publications.666https://www.audeering.com/research/opensmile/ We focus on a subset of the eGEMAPS feature set (Eyben et al., 2015); in particular, we extract the fundamental frequency (F0), the first formant (F1), and the first four Mel Frequency Cepstral Coefficients (MFCCs 1-4). Each features was chosen for its relevance to sociolinguistic phenomena.
F0 is related to vocal pitch, and captures the lowest frequency of oscillations of a voiced sound. In addition to varying across speakers, speakers manipulate this frequency in speaking, such as to mark emphasis or questions. Formants capture additional resonances or harmonics, with F1 being commonly associated with vowel pronunciation. It has been further employed to describe variation in speakers’ dialect, sex, and age Hagiwara (1997); Kent and Vorperian (2018). Finally, MFCCs are features designed to capture the short-term power spectrum of sounds, and can be used in concert with other features to measure general vocal characteristics for a number of downstream tasks such as emotion recognition Luengo et al. (2005). openSMILE measures these features at a high frequency which incurs a large storage requirement; therefore, we collapse these to the token-level, keeping the mean of each feature over each word’s duration.
@@ 3.4 Identifying Speakers
A key limitation of Whisper is that it does not distinguish between speakers. In order to measure turn-level information from podcast conversations, we attempt to match speaking turns to speakers, using a combination of off-the-shelf and newly developed tools.
Speaker Diarization:
@@
As a first step, we use pyannote (Bredin, 2023) to split audio files into individual speaker turns, using a process known as diarization.777Due to computational limitations and the speed of diarization, we only diarize 370K episodes in our initial dataset release, sampled randomly from the full set of transcribed episodes, but will update this in subsequent data releases. pyannote determines the number of distinct speakers in a conversation, matching these speakers to generic labels (e.g., speaker1, speaker2, etc.). Each audio segment is then assigned to one of these speakers, which allows us to map each token from transcripts to a corresponding speaker turn, using audio timing information. Validation against a sample of professional transcripts reveals a diarization word error rate of 2.1% on average for a small sample of two-person conversations. (See Appendix B for details). After excluding speakers that make up less than 5% of the total speaking time in an episode, we find that 37% of episodes in our diarized data have only a single speaker, with 39% having two speakers, and the rest having 3.
Identifying Host and Guest Names
@@
Unfortunately, mapping generic speaker IDs to named speakers or speaker roles is difficult, as this information is not reliably included in episode metadata in a consistent format. As a step towards this, we develop a model to identify the names and voices of the host(s) and guest(s) for each episode.
@@
After preliminary exploration, we determined that transcripts were the most reliable sources of this information, as hosts and guests are typically introduced by name (e.g., “I am your host …”, etc.).888In addition, some hosts never introduce themselves, or provide only their first name, and thus remain effectively anonymous in our data. To determine speaker names and match them to voices, we develop a pipeline to first identify candidate names in text, and then classify each of these names as Host, Guest, or Neither, with the third category representing people mentioned but not appearing in the episode. To make these assignments, we rely on the direct and indirect linguistic cues surrounding mentions of a speaker’s name that indicate their role. We first use spaCy Honnibal et al. (2020) to identify all named Person entities that occur in an episode description, or in the first 350 words of the transcript, filtering this list to keep only mentions that include a first and last name.
@@
To generate training data for this task, we collect three human judgments from Profilic annotators on a subset of 2,000 candidate entities in context. Annotators label one target entity at a time, see the podcast description, episode description, and 300 words of the transcript, and provide a judgment of Host, Guest, or Neither. Despite some ambiguous cases, annotators had relatively high chance-corrected agreement on this task (Krippendorff’s =0.77). Final labels were obtained by aggregating annotations using MACE Hovy et al. (2013), resulting in labels for 858 Hosts, 639 Guests, and 503 Neither. Using these annotations, we fine tune a RoBERTa model to classify a named entity based on its mean-pooled embeddings, which achieves 0.87 cross-validation accuracy and 0.88 accuracy on a held out test set. We then apply this model to all episodes for which there is at least one identified candidate entity. For diarized podcasts where our model identifies exactly one of the speakers as the Host, we heuristically map that person to the first voice which says the host’s name. For additional details on speaker labeling, see Appendix C.
@@ Results
Using our role annotation model, we identify participants in all episodes; over 550K episodes (49%) of the dataset have at least one identified Host or Guest, including 386K Host names and 535K Guest names. Based on these inferred participants, we find that the plurality of podcasts have a single Host, followed by a single Host and single Guest. (See Figure 8 in the Appendix). During manual review, we found a few notable false positives, such as Ryan Reynolds, who appears in a common podcast ad in which he is labeled as a Host. Nevertheless, we verify that some guests truly are prolific, such as Matt Ridley, who appears as a guest on 21 episodes in our data. Further refinements could improve on our host and guest identification by combining information across episodes and using external knowledge bases, but these labels nevertheless permit further investigation into how guests connect shows across categories (§4), and we leave such extensions of our model as future work.
@@ 3.5 Summary of Our Dataset
Altogether, our data includes rich multimodal information spanning the episode and speaker-turn levels. At the episode level, we provide transcript text alongside category, duration, publication date, and inferred host and guest information. At the turn level, we provide transcripts split into segments of speech corresponding to different speakers, start and end timestamps, averaged audio features for each segment, and their inferred host and guest roles where possible. We release episode-level data for the full set of 1.1 million podcast episodes and speaker turn data for the subset of 370K diarized episodes (see Appendix E for details).